# Homotopy Continuation Made Easy: Regression-based Online Simulation of Starting Problem-Solution Pairs

### v1

@@ Abstract
While automatically generated polynomial elimination templates have sparked great progress in the field of 3D computer vision, there remain many problems for which the degree of the constraints or the number of unknowns leads to intractability. In recent years, homotopy continuation has been introduced as a plausible alternative. However, the method currently depends on expensive parallel tracking of all possible solutions in the complex domain, or a classification network for starting problem-solution pairs trained over a limited set of real-world examples. Our innovation consists of employing a regression network trained in simulation to directly predict a solution from input correspondences, followed by an online simulator that invents a consistent problem-solution pair. Subsequently, homotopy continuation is applied to track that single solution back to the original problem. We apply this elegant combination to generalized camera resectioning, and also introduce a new solution to the challenging generalized relative pose and scale problem. As demonstrated, the proposed method successfully compensates the raw error committed by the regressor alone, and leads to state-of-the-art efficiency and success rates while running on CPU resources, only.
@@ 1 Introduction
While polynomial elimination techniques have lead to significant progress in the solution of geometric vision problems, the methods are limited as the complexity of the template search grows uncontrollably with the degree and dimensionality of the problem at hand. Furthermore, owing to their failure to respect inequality constraints and rule out practically infeasible solutions, high degrees or dimensionalities naturally lead to an elevated number of algebraically possible solutions and—assuming that it can be found at all—large elimination templates [44, 64, 63, 13, 14, 62]. This in turn causes problems in terms of numerical stability [62]. In recent years, a different solution strategy has therefore gained popularity: Homotopy Continuation (HC) [60, 4, 12]. The method proceeds by starting from a known problem-solution pair, and then tracks its roots to the target problem at hand during a step-wise interpolation of polynomial coefficients. The method has been used to successfully solve challenging tri-focal relative pose problems [17]. While promising, HC remains expensive as it generally requires the parallel continuation of many roots in the complex domain in order to identify all real solutions to a target problem, followed up by disambiguation.
@@
  • We do not train a classifier on a fixed set of problem-solution pairs taken from a real-world dataset. Instead, we propose to employ a solution regression network trained over an arbitrarily large set of simulated input correspondences for the considered geometric problem. The network is hence not limited by a finite set of candidate problems, and aims at being a general solution approximator for any instance of the polynomial problem.
@@ 3 Polynomial Solver: Homotopy Continuation
The remainder of this paper—our key contribution—addresses the second question of how to find a good starting system. Traditional methods (e.g. Ab initial [17], monodromy [12, 6]) rely on algebraic geometry theories developed for coefficient-parameter homotopies, and tracking all roots from a small set of fixed, known problem-solution pairs in the complex domain may often enable a solution [55]. However, tracking all the isolated paths is expensive. A recent study proposes the use of a classifier to pick a good starting pair with as high as possible success probability upon tracking only a single root [24], which is much faster. Similarly, in this work, we propose a novel geometry-informed strategy that combines a neural regressor with an online simulator to produce a problem-solution pair and then track a single root by HC. Even on challenging problems for which no traditional solver has been found, it can find the solution with a success rate of over $95\%$.
@@
Since being introduced in 1980s, a variety of homotopies have been developed to serve the same purpose: how to solve the problem exactly as expected. The representative one would be coefficient-parameter homotopy. Its formulation provides the insight that if we can find all the solutions to a general coefficient-parameter polynomial, then we can find all the other members of that polynomial family. In fact, as pointed out in [55], in engineering problems (e.g. geometry vision problems) we are dealing with natural coefficient-parameter polynomials, that is coefficients of polynomials derived from visual features. Since the scope of this paper lies on finding a good starting solution that is not highly depend the path tracking, we kindly refer the reader to [55] for an in-depth discussion of other homotopy types. As shown in the later applications, straight segment homotopies provide near-perfect success rates when working with our simulator. We also want to note that all the other homotopy types could be equally applied in conjunction with our simulator.
@@ 4 Find a Good Starting Solution: Regression Network
Following recent developments in deep learning, multiple attempts have been made to for example learn camera pose using a neural network [25, 48, 52]. The first two methods extract features from images and points and use them in a regression network to predict the pose. The third method is less mature and aims at predicting pose directly from geometric correspondences, which is also the aim of the present work. The bottleneck in regressing pose from correspondences arises from two concerns: 1) For simple geometric problems, we often have traditional methods that are efficient and accurate, and 2) For hard geometric problems, regression networks demonstrate a poor ability to predict accurate solutions. Our key insight is that while the predicted pose may not be sufficiently accurate, it can still serve as a good starting solution for HC.
@@ 5 Find a Consistent Starting Problem: Online Simulator
6 Toy example: Generalized Perspective-n-Points
Our first concrete, practical application targets generalized camera pose estimation from 2D-3D correspondences. We use the Unified Perspective-n-Point (UPnP) algorithm [30] which formulates a polynomial system corresponding to the first-order optimality conditions of the sum of squared object space distances. The method can be used to solve both central and non-central camera resectioning problems from an arbitrary number of points (including minimal). It can be efficiently solved by the Gröbner basis method since—after variable elimination—it only involves $4$ unknowns. In this section, we solve UPnP by our newly proposed method and compare it against conventional approaches, a raw regression network, as well as a parallel HC implementation tracking all roots.
@@ 6.2 Regressor and online simulator
where $N$ is the number of correspondences used, and the output is the quaternion vector. The network consists of $4$ 1D convolutional layers and a fully connected layer at the end to convert the output to a quaternion vector with the unit norm. Given the rotation predicted by the network, the translation can be recovered from correspondences and estimated rotation. We refer the reader to [30] for the details. For the online simulator, we can get reprojected $\hat{\mathbf{f}}_{i}$ from (5). Algorithm 1 summarizes all stages, including the final HC over (6). Note that in line 6 of Algorithm 1, the starting problem $\mathbf{G}(\mathbf{x})$ is constructed by replacing the coefficient matrix $\mathbf{M}$ computed by $(\mathbf{p}_{i},\mathbf{f}_{i},\mathbf{v}_{i})$ in (6) with $\hat{\mathbf{M}}$ computed by $(\mathbf{p}_{i},\hat{\mathbf{f}}_{i},\mathbf{v}_{i})$.
@@ 6.3 Experiments
Table 1: Performance comparison of simulator HC, a traditional method, the regressor alone and parallel exhaustive tracking HC on noisy synthetic data. All results are averaged over $1000$ trials. The success rate is computed for rotation errors with a threshold of $2$ degree.
Method Succ. $\mathcal{E}_{\mathbf{R}}(deg)$ $\mathcal{E}_{\mathbf{t}}(\%)$ Time(msec)
 Gröbner Basis 100.0% 0.031 0.489 1.254
  Only Net 32% 7.204 32.162 0.439
 CPU HC 99.9% 0.032 0.499 9.796
 GPU HC 99.9% 0.032 0.499 1.002
  LM 53.6% 28.721 2255.9 32.580
 Simulator HC 99.9% 0.137 0.737 0.160
@@
Data. We randomly generate $4$ cameras in a cube $[-1,1]^{3}$ and randomly sample $16$ 3D points in the box $[-1,1]^{2}\times[4,8]$. 16 samples would not be a good choice for a ransac solver, but in the present context the idea is to set a moderate and notably equal challenge to all compared solvers. Furthermore, for UPnP the number of constraints appearing in the polynomial system remains unchanged as it analyses first-order optimality rather than individual incidence relations. Elements of the rotation angles are randomly generated in the interval $[-\pi/2,\pi/2]$, respectively. And the translation is randomly generated in the unit cube. We then generate the correspondences from rays that point to a 3D point. To add noise, we first extract an orthogonal plane from each ray direction vector and add random noise in pixels by assuming a focal length of 800. The perturbed ray vectors are renormalized. The noise level is set to $2$ pixels. To train the pose regression network, we simulate a dataset of $2400$ samples. We use a 2.70GHz Intel(R) Xeon(R) CPU for evaluation and an NVIDIA GeForce RTX 4090 for network training.
@@
Methods. For HC, the stepsize $\delta t$ and the maximum number of iterations of Newton’s method are set to $0.02$ and $5$, respectively. We compare the Gröbner Basis solver [30] and the popular local method Levenberg Marquardt (LM) [1]. We also test the state-of-the-art parallel HC solver developed by [8, 7], and use both its CPU and GPU [^2] Note that GPU warmup overhead is ignored when averaging over multiple runs. version. Given that parallel HC tracks all solutions, we manually pick the one with the smallest error. Note that the polynomial system from (6) is overdetermined as it has 4 unknowns and 5 equations, and CPU/GPU-HC is only designed to solve square systems. Hence, we use a modified version of UPnP with Cayley parameters in CPU/GPU-HC to track $27$ solutions found by monodromy [12, 6]. This has no impact on the success rate, and will at best benefit this algorithm in terms of runtime owing to smaller Jacobian matrices.
@@
Results. The rotation error is measured by $\mathcal{E}_{\mathbf{R}}=2\,\arcsin(\|\hat{\mathbf{R}}-\mathbf{R}_{gt}\|/2\sqrt{2})$ in degrees as suggested by [11] and the translation error is measured by the $\mathcal{E}_{\mathbf{t}}=\|\hat{\mathbf{t}}-\mathbf{t}_{gt}\|/\|\mathbf{t}_{gt}\|\times 100$ in percentage. Table 1 presents the comparison between the traditional method, the regressor alone, parallel exhaustive tracking HC, and our simulator HC. We present the success rate as the percentage of rotation errors smaller than $2$ degrees. It can be observed that on this simple task, except for the local method, all solvers succeed in over $99.99\%$ of trials. Meanwhile, online simulator HC benefits from the single-track strategy and turns out to be the fastest one. Although the pose learnt by the regressor is not accurate enough, it does turn out to be a good start solution for our simulator HC.
@@ 7 Generalized Relative Pose and Scale
Our final application targets a hard minimal problem: Generalized Relative Pose and Scale (GRPS) estimation from 2D-2D correspondences [31]. The GRPS problem is a natural extension of the generalized relative pose problem and arises in point-less registration of partial view-graphs in structure-from-motion. Owing to the fact that SfM results are generally up to scale, an inherent scale factor needs to be resolved. Such problem was first introduced by Kneip et al. [31]. It has 140 solutions if a unique rotation representation is used, and can be solved by multi-start optimization to overcome local minima. Later on, a traditional polynomial solver based on the Gröbner basis method has been proposed [35, 39]. With a $144\times 284$ elimination template, the solver finds all $140$ rotations within reasonable time. In this section, we solve GRPS by our proposed method and compare it against the existing approach, a raw regression network, as well as a parallel HC implementation tracking all roots.
@@ 7.3 Simulation Experiments
Methods. The stepsize for homotopy continuation is set to $0.05$ in this scenario. We compare against the Eigen solver [31], Gröbner Bases solver [39] and the CPU and GPU versions of parallel HC which tracks $140$ solutions found by Monomdromy [6]. Note that to measure numerical errors, we just pick the best solution with the smallest error. In practice, traditional HC needs to evaluate additional information to select a single solution (e.g. euclidean distance). For a complicated problem like GRPS, evaluating $140$ solutions would still induce non-neglectable time-consumption.
The pose regression network is trained over $64000$ samples.
Metric. We measure the rotation error in degrees and the relative translation and scale error in percentage using
$\mathcal{E}_{\mathbf{R}}$ $\;=2\arcsin(\|\hat{\mathbf{R}}-\mathbf{R}_{gt}\|/2\sqrt{2})$ (12)
$\mathcal{E}_{\mathbf{t}}$ $\;=\|\hat{\mathbf{t}}-\mathbf{t}_{gt}\|/\|\mathbf{t}_{gt}\|\times 100$
$\mathcal{E}_{s}$ $\;=|\hat{s}-s_{gt}|/s_{gt}\times 100.$
@@
Figure 4: Experimental results on GRPS. Fig. 4(a) Error distribution over $1000$ trials on noise-free data. The camera number is set to be $3$. Considering the rotation error, except for Gröbner Bases, the other minimal solvers all present a similar error distribution with around $70\%$ success rate, and the eigen solver presents only around $60\%$ success rate. The proposed $8$pts simulator HC presents $96.3\%$ success rate where the Gröbner Bases has $100\%$ success rate. Fig. 4(b) The boxplot of running time comparison for the proposed simulator HC and other methods. As expected, on average the proposed online simulator HC running on CPU is the fastest. It is about 6×6$\times$6 × faster than GPU-HC since we only track one solution, 20×20$\times$20 × faster than Gröbner Bases and about 90×90$\times$90 × faster than CPU-HC. Fig. 4(c) Error statistics of simulator HC with respect to different noise levels ranging from $0.2$ pixel to $1.0$ pixel.
Time Comparison. The main computational benefit of the proposed online simulator HC comes from single solution-tracking, which can be effectuated efficiently even on the CPU. Figure 4(b) demonstrates the efficiency gain of our $8$pts simulator HC over the eigen solver and parallel HC variants. As expected, by tracking only one solution curve, we are significantly faster than parallel exhaustive tracking HC, even if the latter runs on GPU. The eigen solver is an optimization-based solver with high complexity and ends up being slowest. The Gröbner Bases solver dealing with $144\times 284$ elimination template still costs a considerable time, even for just computing solutions for rotation, and not translation or scale. Therefore, in the following, we only present $8$-points simulator HC results.
@@
We would like to note that Gröbner Bases is solving a different formulation from Eq. 9. For the sake of simplifying the problem to make it easy to solve by an elimination template, it eliminates translation and scale [56]. But the resulting coefficient matrix is $130\times 275$ and still needs a large number of operations. Although the formulation of Eq. 9 with eliminated depths faces a slightly lower success rate, its number of operations is much smaller than Gröbner Bases. On the other hand, the proposed 8pt simulator HC provides a comparably high success rate as Gröbner Bases and is still 20×20$\times$20 × faster.
Table 2: RANSAC experiments for the GRPS problem. Image points are randomly perturbed by 2-pixel uniform noise. The outlier ratio is gradually increased from $10\%$ to $50\%$. We use vanilla RANSAC with at most $200$ iterations. As can be observed, outlier ratios of up to $40\%$ are easily handled within 200 iterations, and the success rate remains above $50\%$ even for $50\%$ outliers.
Ratio Succ. (%) $\mathcal{E}_{\mathbf{R}}(deg)$ $\mathcal{E}_{\mathbf{t}}(\%)$ $\mathcal{E}_{\mathbf{s}}(\%)$ Iters Time(sec)
10% 100.0 0.09 2.03 1.98 125 0.848
20% 100.0 0.10 2.11 2.23 193 1.305
30% 99.7 0.14 3.40 3.17 200 1.353
40% 91.1 0.29 7.57 7.81 200 1.347
50% 53.0 0.56 28.34 26.33 200 1.348
@@
Evaluation inside RANSAC. To conclude, we embed simulator HC into RANSAC to test its practical usefulness. We synthesize $200$ correspondences between $5$ cameras, and add outliers by replacing part of the image points in one of the views by the projection of different, arbitrary 3D points. Uniformly sampled pixel noise from $[-2,2]$ is added to both coordinates of the image points by again assuming a virtual focal length of 800 px. We terminate RANSAC if it reaches $99\%$ confidence or $200$ iterations, and the inlier threshold is set to $0.01$ for the distance measured by the normalized reprojection error. Table 2 shows the results. The success rate is measured by a rotation error threshold of $2$ degrees and the indicated errors are the average accuracy of successful cases. Overall, simulator HC embedded into RANSAC can easily handle up to $40\%$ outliers, which suggests high practicality.
@@
seq. $\mathcal{E}_{\mathbf{R}}$(deg) $\mathcal{E}_{s}(\%)$ $\mathcal{E}_{\mathbf{t}_{\text{ang}}}$(deg) #iters time(sec) #inliers Succ. (%)
CAB 1.23 47.76 12.086 388 2.59 137 63
HGE 0.65 47.93 10.27 167 1.11 242 78
LIN 0.59 39.27 5.32 88 0.59 298 82

### v2

# Simulator HC: Regression-based Online Simulation of Starting Problem-Solution Pairs for Homotopy Continuation in Geometric Vision

@@ Abstract
While automatically generated polynomial elimination templates have sparked great progress in the field of 3D computer vision, there remain many problems for which the degree of the constraints or the number of unknowns leads to intractability. In recent years, homotopy continuation has been introduced as a plausible alternative. However, the method currently depends on expensive parallel tracking of all possible solutions in the complex domain, or a classification network for starting problem-solution pairs trained over a limited set of real-world examples. Our innovation lies in a novel approach to finding solution-problem pairs, where we only need to predict a rough initial solution, with the corresponding problem generated by an online simulator. Subsequently, homotopy continuation is applied to track that single solution back to the original problem. We apply this elegant combination to generalized camera resectioning, and also introduce a new solution to the challenging generalized relative pose and scale problem. As demonstrated, the proposed method successfully compensates the raw error committed by the regressor alone, and leads to state-of-the-art efficiency and success rates.
@@ 1 Introduction
While polynomial elimination techniques have led to significant progress in the solution of geometric vision problems, the methods are limited as the complexity of the template search grows uncontrollably with the degree and dimensionality of the problem at hand. Furthermore, owing to their failure to respect inequality constraints and rule out practically infeasible solutions, high degrees or dimensionalities naturally lead to an elevated number of algebraically possible solutions and—assuming that it can be found at all—large elimination templates [45, 65, 64, 13, 14, 63]. This in turn causes problems in terms of numerical stability [63]. In recent years, a different solution strategy has therefore gained popularity: Homotopy Continuation (HC) [61, 4, 12]. The method proceeds by starting from a known problem-solution pair, and then tracks its roots to the target problem at hand during a step-wise interpolation of polynomial coefficients. The method has been used to successfully solve challenging tri-focal relative pose problems [17]. While promising, HC remains expensive as it generally requires the parallel continuation of many roots in the complex domain in order to identify all real solutions to a target problem, followed up by disambiguation.
@@
  • We do not train a classifier on a fixed set of problem-solution pairs taken from a real-world dataset. Instead, we propose to employ a solution regression network trained over an arbitrarily large set of simulated input correspondences for the considered geometric problem. The network is hence not limited by a finite set of candidate problems, and aims at being a general solution approximator for any instance of the polynomial problem. What’s more, learning a solution is much easier than learning a proper problem-solution pair.
@@ 3 Polynomial Solver: Homotopy Continuation
The remainder of this paper—our key contribution—addresses the second question of how to find a good starting system. Traditional methods (e.g. Ab initial [17], monodromy [12, 6]) rely on algebraic geometry theories developed for coefficient-parameter homotopies, and tracking all roots from a small set of fixed, known problem-solution pairs in the complex domain may often enable a solution [56]. However, tracking all the isolated paths is expensive and in practice multiple solutions cost more time to select the best. A recent study proposes the use of a classifier to pick a good starting pair with as high as possible success probability upon tracking only a single root [25], which is much faster. Similarly, in this work, we propose a novel geometry-informed strategy that combines a neural regressor with an online simulator to produce a problem-solution pair and then track a single root by HC. Even on challenging problems for which no traditional solver has been found, it can find the solution with a success rate of over $95\%$.
@@
Since being introduced in 1980s, a variety of homotopies have been developed to serve the same purpose: how to solve the problem exactly as expected. The representative one would be coefficient-parameter homotopy. Its formulation provides the insight that if we can find all the solutions to a general coefficient-parameter polynomial, then we can find all the other members of that polynomial family. In fact, as pointed out in [56], in engineering problems (e.g. geometry vision problems) we are dealing with natural coefficient-parameter polynomials, that is coefficients of polynomials derived from visual features. Since the scope of this paper lies on a novel starting problem-solution pair finding and it is not highly dependent on the path tracking, we kindly refer the reader to [56] for an in-depth discussion of other homotopy types. As shown in the later applications, straight segment homotopies provide near-perfect success rates when working with our simulator. We also want to note that all the other homotopy types could be equally applied in conjunction with our simulator.
@@ 4 Find a Good Starting Solution: Regression Network
Following recent developments in deep learning, multiple attempts have been made to, for example, learn camera pose using a neural network [26, 49, 53]. The first two methods extract features from images and points and use them in a regression network to predict the pose. The third method is a more preliminary approach and aims at predicting pose directly from geometric correspondences, which is also the aim of the present work. The bottleneck in regressing pose from correspondences arises from two concerns: 1) For simple geometric problems, we often have traditional methods that are efficient and accurate, and 2) For hard geometric problems, regression networks demonstrate a poor ability to predict accurate solutions. Our key insight is that while the predicted pose may not be sufficiently accurate, it can still serve as a good starting solution for HC.
@@ 5 Find a Consistent Starting Problem: Online Simulator
6 Generalized Perspective-n-Points
Our first concrete, practical application targets generalized camera pose estimation from 2D-3D correspondences. We use the Unified Perspective-n-Point (UPnP) algorithm [31] which formulates a polynomial system corresponding to the first-order optimality conditions of the sum of squared object space distances. The method can be used to solve both central and non-central camera resectioning problems from an arbitrary number of points (including minimal). It can be efficiently solved by the Gröbner Bases method since—after variable elimination—it only involves $4$ unknowns. In this section, we solve UPnP by our newly proposed method and compare it against conventional approaches, a raw regression network, as well as a parallel HC implementation tracking all roots.
@@ 6.2 Regressor and online simulator
where $N$ is the number of correspondences used, and the output is the quaternion vector. The network consists of $4$ 1D convolutional layers and a fully connected layer at the end to convert the output to a quaternion vector with the unit norm. Given the rotation predicted by the network, the translation can be recovered from correspondences and estimated rotation. We refer the reader to [31] for the details. For the online simulator, we can get reprojected $\hat{\mathbf{f}}_{i}$ from (5). Algorithm 1 summarizes all stages, including the final HC over (6). Note that in line 5 of Algorithm 1, the starting problem $\mathbf{G}(\mathbf{x})$ is constructed by replacing the coefficient matrix $\mathbf{M}$ computed by $(\mathbf{p}_{i},\mathbf{f}_{i},\mathbf{v}_{i})$ in (6) with $\hat{\mathbf{M}}$ computed by $(\mathbf{p}_{i},\hat{\mathbf{f}}_{i},\mathbf{v}_{i})$.
@@ 6.3 Experiments
Data. We randomly generate $4$ cameras in a cube $[-1,1]^{3}$ and randomly sample $16$ 3D points in the box $[-1,1]^{2}\times[4,8]$. 16 samples would not be a good choice for a RANSAC solver [18], but in the present context the idea is to set a moderate and notably equal challenge to all compared solvers. Furthermore, for UPnP the number of constraints appearing in the polynomial system remains unchanged as it analyses first-order optimality rather than individual incidence relations. Elements of the rotation angles are randomly generated in the interval $[-\pi/2,\pi/2]$, respectively, and the translation is randomly generated in the unit cube. We then generate the correspondences from rays that point to a 3D point. The noise level is set to $2$ pixels. To train the pose regression network, we simulate a dataset of $2400$ samples, which is sufficient for reliable Simulator HC initialization. We use a 2.70GHz Intel(R) Xeon(R) CPU and an NVIDIA GeForce RTX 4090 for network training and evaluation.
@@
Methods. For HC, the stepsize $\delta t$ and the maximum number of iterations of Newton’s method are set to $0.02$ and $5$, respectively. We compare the Gröbner Bases solver [31] and the popular local method Levenberg Marquardt (LM) [1]. We also test the state-of-the-art parallel HC solver developed by [8, 7], and use both its CPU and GPU [^2] Note that GPU warmup overhead is ignored when averaging over multiple runs. version. Given that parallel HC tracks all solutions, we manually pick the one with the smallest error.
@@
Results. The rotation error is measured by $\mathcal{E}_{\mathbf{R}}=2\,\arcsin(\|\hat{\mathbf{R}}-\mathbf{R}_{gt}\|/(2\sqrt{2}))$ in degrees as suggested by [11] and the translation error is measured by the $\mathcal{E}_{\mathbf{t}}=\|\hat{\mathbf{t}}-\mathbf{t}_{gt}\|/\|\mathbf{t}_{gt}\|\times 100$ in percentage. Table 1 compares the traditional method, parallel exhaustive tracking HC, and different types of initial solutions for our simulator HC. We present the success rate as the percentage of rotation errors smaller than $2$ degrees. It can be observed that on this simple task, except for the local method, traditional methods, as well as simulator HC using predicted initials, succeed in $100\%$ of trials. Meanwhile, online simulator HC benefits from the single-track strategy and turns out to be the fastest one. We also perform an ablation study on the proposed simulator HC, using both predicted and random initializations. Although the pose learned by the regressor is not highly accurate and only slightly outperforms random poses, it serves as a reasonable starting solution for our simulator HC. In contrast, LM as a local method is unable to leverage the potential of the predicted initializations, as the simulator HC does.
Table 1: Performance comparison of simulator HC, a traditional method, and parallel exhaustive tracking HC on noisy synthetic data. All results are averaged over $1000$ trials. The success rate is computed for rotation errors with a threshold of $2$ degree.
Method Succ. (%) $\mathcal{E}_{\mathbf{R}}(deg)$ $\mathcal{E}_{\mathbf{t}}(\%)$ Time(ms)
Gröbner Bases 100 0.031 0.489 1.254
CPU HC 100 0.032 0.499 9.796
GPU HC 100 0.032 0.499 1.002
Pred Initial 32 7.204 32 0.031
Pred Initial + LM 54 28.72 2256 0.528
Rand Initial + Simulator HC 22 91.68 228 0.231
Pred Initial + Simulator HC 100 0.032 0.513 0.316
@@ 7 Generalized Relative Pose and Scale
Our final application targets a hard minimal problem: Generalized Relative Pose and Scale (GRPS) estimation from 2D-2D correspondences [32]. The GRPS problem is a natural extension of the generalized relative pose problem and arises in point-less registration of partial view-graphs in structure-from-motion. Owing to the fact that SfM results are generally up to scale, an inherent scale factor needs to be resolved. Such problem was first introduced by Kneip et al. [32]. It has 140 solutions if a unique rotation representation is used, and can be solved by multi-start optimization to overcome local minima. Later on, a traditional polynomial solver based on the Gröbner Bases method has been proposed [36, 40]. With a $144\times 284$ elimination template, the solver finds all $140$ rotations within reasonable time. In this section, we solve GRPS by our proposed method and compare it against the existing approach, as well as a parallel HC implementation tracking all roots.
@@ 7.3 Simulation Experiments
Methods. The stepsize for homotopy continuation is set to $0.05$ in this scenario. We compare against the Eigen solver [32], Gröbner Bases solver [40] and the CPU and GPU versions of parallel HC which tracks $140$ solutions found by Monodromy [6]. Note that to measure numerical errors, we just pick the best solution with the smallest error. In practice, traditional HC needs to evaluate additional information to select a single solution (e.g. euclidean distance). For a complicated problem like GRPS, evaluating $140$ solutions would still induce non-neglectable time-consumption. Since GRPS is a hard problem, we train the regressor using a larger simulated dataset of size $64000$.
Metric. We measure the rotation error in degrees and the relative translation as in Sec. 6.3 and scale error in percentage using
$\mathcal{E}_{s}=|\hat{s}-s_{gt}|/s_{gt}\times 100.$ (12)
@@
Figure 4: Experimental results on GRPS. Fig. 4(a) Error distribution over $1000$ trials on noise-free data. The camera number is set to be $3$. Considering the rotation error, except for Gröbner Bases, the other minimal solvers all present a similar error distribution with around $70\%$ success rate, and the eigen solver presents only around $60\%$ success rate. The proposed $8$pts simulator HC presents $96.3\%$ success rate where the Gröbner Bases has $100\%$ success rate. Fig. 4(b) The boxplot of running time comparison for the proposed simulator HC and other methods, where the number represents the median time. As expected, on average the proposed online simulator HC running on CPU is the fastest. It is about 5×5$\times$5 × faster than GPU-HC since we only track one solution, 17×17$\times$17 × faster than Gröbner Bases and about 40×40$\times$40 × faster than CPU-HC. Fig. 4(c) Error statistics of simulator HC with respect to different noise levels ranging from $0.2$ pixel to $1.0$ pixel.
Time Comparison. The main computational benefit of the proposed online simulator HC comes from single solution-tracking, which can be effectuated efficiently even on the CPU. Figure 4(b) demonstrates the efficiency gain of our $8$pts simulator HC over the eigen solver and parallel HC variants. As expected, by tracking only one solution curve, we are significantly faster than parallel exhaustive tracking HC, even if the latter runs on GPU. The eigen solver is an optimization-based solver with high complexity and ends up being slowest. The Gröbner Bases solver dealing with the large elimination template still costs a considerable time, even for just computing solutions for rotation, and not translation or scale.
@@
We would like to note that Gröbner Bases is solving a different formulation from Eq. 9. In order to make it easy to be solved by a smaller elimination template, it eliminates translation and scale [57]. But the resulting coefficient matrix is $130\times 275$ and still needs a large number of operations. Although the formulation of Eq. 9 with eliminated depths faces a slightly lower success rate, its number of operations is much smaller than Gröbner Bases. On the other hand, the proposed 8pt simulator HC provides a comparably high success rate as Gröbner Bases and is still 17×17$\times$17 × faster.
Table 2: RANSAC experiments for the GRPS problem using Gröbner Bases $\mid$ simulator HC. Image points are randomly perturbed by 2-pixel uniform noise. We use vanilla RANSAC with at most $200$ iterations. As can be observed, outlier ratios of up to $40\%$ are easily handled within 200 iterations. And regressor-based simulator HC is consistently faster than Gröbner Bases.
Ratio Succ. (%) $\mathcal{E}_{\mathbf{R}}$(deg) # Iters Time(s)
10%   99.5 $\mid$ 100 0.10 $\mid$ 0.09   93 $\mid$ 125 2.33 $\mid$ 0.85
20%   99.4 $\mid$ 100 0.11 $\mid$ 0.10 171 $\mid$ 193 4.37 $\mid$ 1.31
30%   99.4 $\mid$ 100 0.13 $\mid$ 0.14 200 $\mid$ 200 5.10 $\mid$ 1.35
40% 98.9 $\mid$ 91 0.21 $\mid$ 0.29 200 $\mid$ 200 4.59 $\mid$ 1.35
@@
Evaluation inside RANSAC. To conclude, we embed simulator HC into RANSAC to test its practical usefulness. We synthesize $200$ correspondences between $5$ cameras and replace part of them with outliers. Uniformly sampled pixel noise from $[-2,2]$ is added to both coordinates of the image points by again assuming a virtual focal length of 800 px. We terminate RANSAC if it reaches $99\%$ confidence or $200$ iterations, and the inlier threshold is set to $0.01$ for the distance measured by the normalized reprojection error. Given that Gröbner Bases solver estimates $140$ solutions, we keep the real ones and select the best solution using reprojection errors in parallel to save time for it. Table 2 shows the results. The success rate is measured by a rotation error threshold of $2$ degrees and the indicated errors are the average accuracy of successful cases, results on translation and scale are left in supplementary. Overall, simulator HC embedded into RANSAC can easily handle up to $40\%$ outliers as Gröbner Bases and much faster, which suggests high practicality. Note that Gröbner Bases provides around 46 real solutions each time, which requires a significant amount of additional time for solution selection. Meanwhile, simulator HC does not require this extra step.
@@
seq. $\mathcal{E}_{\mathbf{R}}$(deg) $\mathcal{E}_{s}(\%)$ $\mathcal{E}_{\mathbf{t}}$(deg) #iters Time(s) Succ. (%)
Gröbner Bases CAB 1.18 56.14 5.70 137 2.83 60
HGE 0.62 62.48 4.39 59 1.25 67
LIN 0.60 42.04 3.40 31 0.65 72
simulator HC CAB 1.05 43.20 9.16 310 2.08 68
HGE 0.56 41.52 7.19 134 0.90 84
LIN 0.56 32.19 4.32 66 0.44 86